OpenStudy (anonymous):

: Show that the equation: 5x+9/6x^2-7x-3=2/5 Can be written as: 4x^2-13x-17=0

jhonyy9 (jhonyy9):

25x+45=12x2-14x-6 12x2 -39x -51 =0 /3 4x2 -13x -17 =0

jhonyy9 (jhonyy9):

your answer is right